OpenStudy (anonymous):

Find the Estimated Value 1. 1/8

OpenStudy (anonymous):

Estimated how? Decimal? Integer? I'm not sure what you want (or your teacher wants) here.

OpenStudy (ash2326):

We have \[\frac{1}{8}\] Let's make it easier by multiplying and dividing by 100 \[\frac{1 \times 100}{8\times 100}\] now \[100/8=12.5\] so \[\frac{1}{8}=\frac{12.5}{100}=0.125\]

OpenStudy (anonymous):

Another way to estimate this one: 1/4 is 0.25; 1/8 is half of that, 0.125

OpenStudy (anonymous):

3/4?

OpenStudy (ash2326):

@BrHowk do you know how to divide 30/4?

OpenStudy (anonymous):

@BrHowk are you good at long divison? This tutorial should be really helpful. http://www.youtube.com/watch?v=zyXTr5za8fQ

OpenStudy (anonymous):

BrHawk, you need to estimate 3/4?

OpenStudy (anonymous):

If you have three quarters, how many cents is that?

OpenStudy (anonymous):

@BrHowk after watching that youtube video I sent you, turning fractions like 1/8, 2/3 23/1372 will be a piece of cake. All you do when you turn a fraction into a decimal is divide the top number by the bottom.

OpenStudy (anonymous):

Here's another really good resource. You put in the fraction you want to convert, hit enter, and it will give you a step-by-step solution for the long division that's involved. http://www.webmath.com/fract2dec.html